Cash $ 15,400
  Marketable securities 7,860
  Accounts receivable 12,920
  Inventory 10,650
  Property and equipment 169,500
  Accumulated depreciation (12,100 )
  Total assets $ 204,230
  Liabilities and Stockholders’ Equity
  Accounts payable $ 8,780
  Current notes payable 3,080
  Mortgage payable 4,500
  Bonds payable 21,960
  Common stock 114,600
  Retained earnings 51,310
  Total liabilities and stockholders’ equity $ 204,230


The average number of common stock shares outstanding during 2018 was 870 shares. Net income for the year was $15,000.

Required

Compute each of the following (Round your answers to 2 decimal places.)

a. Current ratio

b. earnings per share

c. quick (acid-test) ratio

d. Return on investment

e. return on equity

f. debt to equity ratio
Answer #1

a) Current ratio = 46830/11860 = 3.95:1

b) Earning per share = 15000/870 = 17.24 per share

c) Quick ratio = 36180/11860 = 3.05:1

d) Return on investment = 15000/204230 = 7.34%

e) Return on equity = 15000/165910 = 9.04%

f) Debt to equity = 38320/165910 = 23.10%